OpenStudy (anonymous):

How do you find the exact values of θ for cscθ= 2 √3 / 3

OpenStudy (phi):

replace cscθ with 1/sinθ so you have \[ \frac{1}{\sin \theta}= \frac{2 \sqrt{3}}{3} \] now flip both sides. what do you get ?

OpenStudy (anonymous):

sinθ= 1/ 2 √3 / 3

OpenStudy (phi):

you flip looks confusing... don't you mean \[ \sin θ = \frac{3}{2\sqrt{3}} \] now divide sqr(3) into 3. what do you get ?

OpenStudy (anonymous):

can you explain the flip...

OpenStudy (anonymous):

the denominator becomes the numerator and the numerator becomes the denominator

OpenStudy (phi):

if you start with \[ \frac{a}{b} \] to flip, you make the top the bottom and vice versa \[\frac{b}{a} \]

OpenStudy (phi):

so in you problem the 2 sqr3 becomes the bottom and 3 becomes the top

OpenStudy (phi):

now the next step is simplify 3/sqr(3)

OpenStudy (anonymous):

Got it. :)

OpenStudy (phi):

you should remember that sqr3 * sqr3 = 3 if you divide both sides by sqr3 you get sqr3 = 3/sqr3 use that to simplify \[ \sin \theta= \frac{3}{2\sqrt{3}} \]

OpenStudy (anonymous):

okay so cscθ = 2√3/3 would be equal to sinθ= 3/ 2√3

OpenStudy (phi):

yes, but you should simplify to \[ \sin\theta= \frac{\sqrt{3}}{2} \] that should look familiar. see http://www.easycalculation.com/trigonometry/trigonometry-tables.php

OpenStudy (anonymous):

can you show me how you got from 3/ 2√3 to √3/2.

OpenStudy (phi):

I did. the idea is \[ 3=\sqrt{3} \cdot \sqrt{3}\] that is the definition of square root if we divide both sides by sqr3 we get \[ \frac{3}{\sqrt{3}}= \sqrt{3} \] that means you can replace \( \frac{3}{\sqrt{3}}\) with \(\sqrt{3} \)

OpenStudy (phi):

another way to simplify is "rationalize" the fraction by multiplying top and bottom by sqr(3) \[ \frac{3}{2\sqrt{3}} \cdot \frac{\sqrt{3}}{\sqrt{3}}= \frac{3\sqrt{3}}{2\cdot 3} = \frac{\cancel{3}\sqrt{3}}{2\cdot \cancel{3}}= \frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

That works better for me. ;) Okay so then I know sin(x)= √3/2 which has an exact value or 60 degrees. If the given domain is -270 degrees<x<180 degrees I can find the other solutions on the unit circle correct?

OpenStudy (anonymous):

another one would be 120 degrees because sin is positive in that quadrant as well.

OpenStudy (phi):

is it -270º ? if so, that is the same as 90º and you want the angle in the 2nd quadrant (90 to 180) the reference angle is 60, and the angle you want is 180-60=120º

OpenStudy (anonymous):

Okay I got the 120 degrees. so what would the exact value be for the - 270 degrees? Wouldn't it be a negative value... so -90 degrees?

OpenStudy (phi):

minus 270 means start at the x-axis and rotate clockwise 270 degrees, you end up at +90 or -270+360 (you can always add 360º) -90 is the same as 270º (numerically: -90+360)

OpenStudy (anonymous):

but how would you state that as an exact value

OpenStudy (phi):

I don't understand the question. I thought it was: if we restrict the domain to -270º to 180º what theta gives csc(theta)=2 √3 / 3

OpenStudy (anonymous):

sorry. This isn't explained very well in the textbook. One of the values is 60 degrees and the other is 120 degrees. I understand that completely, but what about the negative value?

OpenStudy (phi):

if they really mean -270º to +180º then that includes -240º +60º +120º

OpenStudy (anonymous):

-240 because there's the negative value of √3/2 on the unit circle. It's all coming together. I really appreciate your time.

OpenStudy (anonymous):

Now to go one step further and determine the exact values of cos(x) and tan(x) I would use their ratios, right?

OpenStudy (phi):

**-240 because there's the negative value of √3/2 on the unit circle.*** I would not say it like that... it sounds wrong you only want angles A such that sin(A)= +√3/2

OpenStudy (phi):